OpenStudy (anonymous):

Write the equation of a line through the given points in standard form (-6,1), (-1,4) a. 3x - 5y = -23 b. 3x - 5y = -13 c. 3x + 5y = -13 d. 3x + 5y = -23

OpenStudy (dayakar):

equation of the line passing through (x1,y1)=(-6,1) and (x2,y2) =(-1,4) is (y-y1)(x2 -x1) = ((y2- y1) ( x -x1) substite x1, x2, y1,y2 values in the above equation

OpenStudy (anonymous):

i dont undestand

OpenStudy (anonymous):

i don't either

OpenStudy (anonymous):

a and d are the same and so are b and c

OpenStudy (anonymous):

just kidding

OpenStudy (dayakar):

x1= -6, y1= 1 , x2 = -1 ,y2= 4 (y- 1) [-1-1]= [4-1] [x-(-6)] can u simplify it

OpenStudy (anonymous):

thanks for the help you two :)

OpenStudy (dayakar):

yes

OpenStudy (dayakar):

what

OpenStudy (dayakar):

d